BBC Worldwide is setting prices for the Blue-Ray box set of the eleventh season (they would say “series”) of Doctor Who. BBC’s market research department has found that demand for Doctor Who DVDs and Blue-Ray discs appears to be more elastic in the US than in the UK. In particular, the estimated elasticity of demand for the Blue-Ray box set is (Ea=) –1.25 among American consumers, and (Eb=) –1.16 among their British counterparts.

a) Suppose the price of ‘The Complete 11th Series’ in the UK has been set at £45, which at the current exchange rate is $58. What must be the price of the season 11 Blue Ray in the US?
Answer #1

Exchange rate = $58 / 45 = 1.29 dollars per pound

Lerner Index = - 1 / E = (P - MC) / P

For UK, Price = 45

- 1 / Eb = (45 - MC) / 45

- 1 / - 1.16 = (45 - MC) / 45

0.862 = (45 - MC) / 45

38.79 = 45 - MC

MC = 6.21 = 6.21 x ($1.29/) = $8.01

In US,

- 1 / Ea = (P - MC) / P

- 1 / - 1.25 = (P - 8.01) / P

0.8 = (P - 8.01) / P

0.8P = P - 8.01

0.2P = 8.01

P = $40.05